OpenStudy (anonymous):

how do I solve the square root of x+2 -1 (-1 is outside the root) = square root of x-5 ?

OpenStudy (anonymous):

so its \[\sqrt{x+2}-1=\sqrt{x-5}\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

not sure how to square the x+2 part to get the answer

OpenStudy (anonymous):

x=14? I'm not sure because I haven't done this in a while..

OpenStudy (anonymous):

\[\sqrt{x+2}-1=\sqrt{x-5}\] \[\sqrt{x+2}^{2}-1^{2}=\sqrt{x-5}^{2}\] x+2-1=x-5 x=6

OpenStudy (anonymous):

x=14! thanks so much

OpenStudy (anonymous):

question really fast. how did you square the x+2 side?

OpenStudy (anonymous):

x=14? Okay, so what I did was plug in another variable in place of \[\sqrt{x+2}\].. So I made that equal y. Maybe that will be easier, then you solve it regularly, but remember to substitute \[\sqrt{x+2}\] in y.

OpenStudy (anonymous):

oh okay! thanks so much!